OpenStudy (anonymous):

The distance from home plate to the fence in dead center in wrigley field is 400 feet. How far is it from the fence in dead center to third base? How far is it from the fence in dead center to second base?

OpenStudy (campbell_st):

draw a diagram. I've used 90ft as the length of the base paths  1 find the distance from home to 2nd using pythagoras' theorem call it a \[a^2 = 90^2 + 90^2\] \[a=\sqrt{16200} = 90\sqrt{2}\] the distance of 2nd to the fence in dead centre is \[400 - 90\sqrt{?}\] I'll assume the line from home to the fence bisects the right angle at 2nd base  then using the cosine rule, the distance from the fence to 3rd base is \[x^2 = 90^2 + (400 - 90\sqrt{2})^2 - 2 \times 90 \times (400-90\sqrt{2}) \times \cos(135)\] I've done the last part as the 2nd part....